# Modeling Item-Level Dynamic Variability with Residual Diffusion for Bundle Recommendation

### v2

@@
Abstract.
Existing solutions for bundle recommendation(BR) have achieved remarkable effectiveness for predicting the user’s preference for prebuilt bundles. However, bundle-item(B-I) affiliation will vary dynamically in real scenarios. For example, a bundle themed as ‘casual outfit’ may add ‘hat’ or remove ‘watch’ due to factors such as seasonal variations, changes in user preferences or inventory adjustments. Our empirical study demonstrates that the performance of mainstream BR models will fluctuate or even decline regarding item-level variability. This paper makes the first attempt to address the above problem and proposes a novel Residual Diffusion for Bundle Recommendation(RDiffBR) as a model-agnostic generative framework which can assist a BR model in adapting this scenario. During the initial training of the BR model, RDiffBR employs a residual diffusion model to process the item-level bundle embeddings which are generated by BR model to represent bundle theme via a forward-reverse process. In the inference stage, RDiffBR reverses item-level bundle embeddings obtained by the well-trained bundle model under B-I variability scenarios to generate the effective item-level bundle embeddings. In particular, the residual connection in our residual approximator significantly enhances item-level bundle embeddings generation ability of BR models. Experiments on six BR models and four public datasets from different domains show that RDiffBR improves the performance of Recall and NDCG of backbone BR models by up to 23%, while only increased training time about 4%. Codes and datasets are available at https://anonymous.4open.science/r/RDiffBR.
Bundle Recommendation, Diffusion Model, Dynamic Variability Scenario
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; New York,USAisbn: 978-1-4503-XXXX-X/2025/06ccs: Information systems Recommender systems
1. INTRODUCTION
@@
Figure 1. An example shows bundle-item affiliation is changed while bundle theme is relatively stable.
@@
In practice, platforms can add the ‘hat’ to align with summer promotions or remove the ‘watch’ based on user feedback. An intuitive idea is to regard it as a new bundle. However, such item-level changes will not cause bundle theme shifting although bundle theme is jointly determined by the items within the bundle. As shown in Figure 1, the theme of Bundle 2 and Bundle 3 remain ‘casual outfit’ by item variation. It is advisable to give priority to reusing the identification information of the original bundle because this dynamic variability maintains the theme of bundle.
In recent years, lots of works have focused on bundle recommendation and achieved good results. Including factorization models represented by LIRE (Liu et al., 2014), EFM (Cao et al., 2017), attention-based models represented by DAM (Chen et al., 2019b), Attlist (He et al., 2019) and GNN-based models represented by BGCN (Chang et al., 2020), CrossCBR (Ma et al., 2022). However, our experiments have found that the performance of existing mainstream bundle recommendation models will fluctuate or even decline in B-I dynamic variability scenarios we have mentioned above. For instance, we found that on Youshu dataset, Recall@20 of BGCN will decrease by up to 19% as items in bundles increases or decreases(see Section 3.2 for details). This is because existing bundle recommendation models lack the ability to effectively model the item-level dynamic variability when generating bundle representations regarding item-level.
The main contributions of this work are summarized as follows:
$\bullet$ We propose a generative residual diffusion framework for bundle recommendation to address item-level dynamic variability issue.
$\bullet$ Our residual approximator in diffusion can significantly enhance item-level bundle embeddings generation ability of BR models.
$\bullet$ Extensive experiments on four datasets from different domains show that RDiffBR enhances the performance of Recall and NDCG of backbone BR models by up to 23%, while only increased training time about 4%.
Figure 2. Overview of the proposed RDiffBR framework.
2. Methodology
Generally speaking, there is an approache to deal with the item-level dynamic variability. Utilize side information (such as item attributes) to introduce additional features but requires architectural modifications to the original model. In contrast, our RDiffBR does not require additional information. The following will give our technical details.
2.1. Problem Definition
$U(|U|=M)$ is denoted as the user set, $B(|B|=N)$as the bundle set, and $I(|I|=O)$ as the item set. The user-bundle interactions, user-item interactions, and bundle-item affiliations are represented as $X_{M\times N}=\{(u,b)\mid u\in U,b\in B\}$, $Y_{M\times O}=\{(u,i)\mid u\in U,i\in I\}$ and $Z_{N\times O}=\{(b,i)\mid b\in B,i\in I\}$ with a binary value at each entry respectively. The value 1 indicates an observed interaction of user-item or user-bundle, or affiliation of bundle-item. The item-level dynamic variability problem can be formulated as follows:
Output: A well-trained bundle recommendation estimates the probabilty $\hat{y}_{ub}$ of a user $u$ interacting with a bundle $b$ based on $Z^{*}=Z\cup\Delta^{+}\setminus\Delta^{-}$ rather than $Z$.
Here, $\Delta^{+}=\{(b,i)\mid(b,i)\notin Z\}$ and $\Delta^{-}=\{(b,i)\mid(b,i)\in Z\}$ represent dynamic variability bundle-item affiliation. And when an item is replaced by others, this can be treated as the joint operation of $\Delta^{+}$ and $\Delta^{-}$.
2.2. RDiffBR Framework
Considering the successful experience of diffusion as a generative model and the sparsity of input data, our RDiffBR framework introduces a residual diffusion model in the latent space of interactions. As shown in Figure 2, after BR model utilizes an encoder to get bundle-level user/bundle embeddings $E^{BL}_{u},E^{BL}_{b}$ and item-level embeddings user/bundle embeddings $E^{IL}_{u},E^{IL}_{b}$ from $\{X,Y,Z\}$. Since $E^{IL}_{b}$ represents the bundle theme generated by the BR model, we iteratively process $E^{IL}_{b}$ through a forward-backward process, jointly optimizing the BR model and the residual approximator module that we have designed based on the task. In the inference stage, RDiffBR applies this optimized forward-reverse process to $E^{IL}_{b}$ from dynamic variability B-I affiliations $Z^{*}$ to obtain new $E^{IL}_{b}$ for bundle recommendation, enhancing BR model’s ability for this scenario.
2.2.1. Residual Diffusion.
(1) $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\boldsymbol{I}).$
Based on the reparameterization (Ho et al., 2020) and the additivity of independent Gaussian noises (Luo, 2022), we can directly obtain $e_{t}$ from $e_{0}$.
(2) $q(e_{t}|e_{0})=\mathcal{N}(e_{t},\sqrt{\bar{\alpha}_{t}}e_{0},(1-\bar{\alpha_{t}})\boldsymbol{I}),$
where $\bar{\alpha_{t}}=\prod_{t=1}^{t}\alpha_{t},\alpha_{t}=1-\beta_{t},\beta_{t}\in(0,1)$ controls the Gaussian noise scales added at each step $t$. To regulate the added noises in $e_{1:T}$, we follow  (Wang et al., 2023) using a linear variance noise schedule in the forward process:
(3) $1-\bar{\alpha}_{t}=s\cdot\left[\alpha_{\min}+\frac{t-1}{T-1}(\alpha_{\max}-\alpha_{\min})\right],$
where T is the total forward step and s $\in$ (0, 1) controls the noise scale. The condition $\alpha_{min}$ $<$ $\alpha_{max}$in (0, 1) indicates the upper and lower limits of the added noise.
(4) $p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\Sigma_{\theta}(\hat{e}_{t},t)),$
where $\hat{e}_{t}$ is denoised embedding in the reverse step t, $\theta$ and are the learnable parameters of residual approximator. This module are executed iteratively in the reverse process until the reconstruction of final embedding $\hat{e}_{0}$
(5) $q(e_{t-1}|e_{t},e_{0})\propto\mathcal{N}(e_{t-1};\tilde{\mu}(e_{t},e_{0},t),\sigma^{2}(t)\boldsymbol{I}),\mathrm{~{}where}$
$\begin{cases}\tilde{\mu}(e_{t},e_{0},t)=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}e_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}e_{0}\\ \sigma^{2}(t)=\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}},&\end{cases}$
@@ 2.2. RDiffBR Framework
$\tilde{\mu}(e_{t},e_{0},t)$ and $\sigma^{2}(t)I$ are the mean and covariance of $q(e_{t-1}|e_{t},e_{0})$. Following  (Zhao et al., 2024) we can similarity factorize $p_{\theta}(e_{t-1}|e_{t})$:
@@
(6) $p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\tilde{\beta}_{t}\boldsymbol{I})),\quad\mathrm{~{}where}$
@@
(7) $\tilde{e}_{0}=\delta f_{\theta}(\hat{e}_{t},t)+(1-\delta)e_{t^{\prime}},$
where $\tilde{e}_{0}$ is the predicted value of $e_{0}$ by residual approximator with parameter $\theta$. $\delta$ is residual connection weight. $e_{t^{\prime}}$ is noisy embedding in the forward step $t^{\prime}$. Step information t is encoded by sinusoidal positional encoding (Ho et al., 2020), and these two inputs are concated as input.
2.2.2. Loss Function.
@@
(8) $\mathcal{L}_{r}$ $=\mathbb{E}_{q(\mathbf{e}_{t}|\mathbf{e}_{0})}\left[D_{\mathrm{KL}}\left(q\left(\mathbf{e}_{t-1}|\mathbf{e}_{t},\mathbf{e}_{0}\right)\|p_{\theta}\left(\mathbf{e}_{t-1}|\mathbf{e}_{t}\right)\right)\right]$
@@
Furthermore, by substituting Eq.(5) and Eq.(6) into Eq.(8) and simplifying (Wang et al., 2023) we can drive:
(9) $\mathcal{L}_{\mathrm{diff}}=\mathbb{E}_{t\sim\mathcal{U}(1,T)}\mathbb{E}_{q}\left[||e_{0}-\hat{e}_{\theta}(e_{t},e_{t^{\prime}},t)||_{2}^{2}\right],$
@@
(10) $\mathcal{L}=\mathcal{L}_{BR\mathrm{~{}model}}+\lambda\mathcal{L}_{\mathrm{diff}}.$
@@ 12            Take gradient descent step on $\nabla_{\theta}(\mathcal{L})$ to optimize $\theta$;
2.2.3. Inference.
In the inference stage, we take the item-level bundle embedding $e^{*}_{0}$ in the B-I dynamic variability scenario as input. Subsequently, RDiffBR continuously introduces noise into $e^{*}_{0}$ in the sequence $e^{*}_{0}\to e^{*}_{1}\to...\to e^{*}_{T^{\prime}}$ for $T^{\prime}$ steps in the forward process and then we set $\hat{e}_{{T^{\prime}}}=e^{*}_{T^{\prime}}$ to execute reverse process $\hat{e_{T^{\prime}}}\to\hat{e}_{{T^{\prime}}-1}\to...\to\hat{e}_{0}$ to generate new item-level bundle embedding $\hat{e}_{0}$. The reverse ignores the variance and uses $\hat{e}_{t-1}=\mu_{\theta}(\hat{e}_{t},e^{*}_{0},t)$ for directly inference following (Liang et al., 2018). We replace the original $e^{*}_{0}$ with $\hat{e}_{0}$, and then conduct bundle recommendation following the original way of the bundle model. Finally enhancing the ranking of bundles in dynamic variation scenarios. The inference procedure of residual approximator is illustrated in Algorithm 2. Specifically, Line 1 represent the forward process, Lines 2-5 represent the reverse process using a residual approximator to obtain ideal item-level bundle embedding.
@@
3. EXPERIMENTS
Our experiments aim to answer the following research questions:
$\bullet$ RQ1: How does RDiffBR perform on different datasets of backbone bundle models under various degrees of B-I variability?
$\bullet$ RQ2: What is the contributions of residual connection in RDiffBR to its overall performance?
@@ 3. EXPERIMENTS
3.1. Experimental Settings
$\bullet$ Datasets. Four publicly bundle datasets from different domains are summarized in Table 1. Youshu (Chen et al., 2019b) for book list recommendation; NetEase (Cao et al., 2017) for music playlist recommendation; iFashion (Chen et al., 2019a) for fashion outfit recommendation; In addition, we introduce another meal recommendation dataset $\mathrm{MealRec+}$ (Li et al., 2024a), where the meal consisted of individual course items is treated as bundle.
Table 1. Dataset Statistics.
Dataset #U #I #B #U-I #U-B #B-I #Avg.I/B #Dens.B-I
Youshu 8,039 32,770 4,771 138,515 51,377 176,667 37.03 0.11%
NetEase 18,528 123,628 22,864 1,128,065 302,303 1,778,838 77.8 0.06%
iFashion 53,897 42,563 27,694 2,290,645 1,679,708 106,916 3.86 0.01%
$\mathrm{MealRec+}$ 1,575 7,280 3,817 151,148 46,767 11,451 3 0.77%
$\bullet$ Baselines. We implement RDiffBR on six representative BR models with official code released by their authors. BGCN (Chang et al., 2020) utilizes graph convolutional networks to separately learn the representations of bundle view and item view. HyperMBR (Ke et al., 2023a) models two views in hyperbolic space to learn their accurate representations. MIDGN (Zhao et al., 2022) refines the user-item/bundle-item graph of different intents from multiple views. CrossCBR (Ma et al., 2022) improves two separate views by constrastive learning. MultiCBR (Ma et al., 2024a) first fuses the multi-view representations into a unified one, and then adopts a contrastive loss on the unified representations. BundleGT (Wei et al., 2023) learns the strategy-aware user and bundle representation in bundle recommendation. There are still some earlier research outcomes (for instance, BundleNet (Deng et al., 2020)) that have been studied to be less effective than the aforementioned methods. Therefore, we do not take them into consideration.
$\bullet$ Experimental process. According to previous works (Li et al., 2024a) (Ma et al., 2022), the training, validation and testing sets of user-bundle interactions $X$ are randomly splited with specific ratio. It is not feasible to simply add new bundles in bundle-item affiliations $Z$ under the B-I variability assumptions. Therefore, we adjust existing bundle datasets to match this scenario: Let $\{{\mathcal{E}_{1},\mathcal{E}_{2},…,\mathcal{E}_{10}}\}$ denote a uniform partition of $Z$ into ten disjoint subsets. Bundle models are trained and validated on $Z_{\mathrm{train\&val}}=\cup_{i=1}^{5}\mathcal{E}_{i}$. In testing phase, we design $Z_{\mathrm{test}}$ by adjusting $Z_{\mathrm{train\&val}}$ with adding or removing subsets:
(11) $Z_{\mathrm{test}}=\begin{cases}Z_{\mathrm{train\&val}}\cup\left(\mathcal{\cup}_{i=6}^{5+\rho}\mathcal{E}_{i}\right)&\mathrm{if}~{}\rho>0\\ Z_{\mathrm{{train\&val}}}&\mathrm{if~{}\rho=0}\\ Z_{\mathrm{train\&val}}\setminus\left(\mathcal{\cup}_{i=6-|\rho|)}^{5}\mathcal{E}_{i}\right)&\mathrm{if~{}\rho<0}\end{cases}$
Where $\cup$ indicates set union and $\setminus$ indicates set difference. $\rho\in\{-4,-3,-2,-1,0,1,2,3,4,5\}$ reflects level of B-I variation. $\rho>0$ indicates adding set $\Delta^{+}$ to $Z_{\mathrm{test}}$, $\rho<0$ indicates removing set $\Delta^{-}$ from $Z_{\mathrm{test}}$ and $\rho=0$ indicates $Z_{\mathrm{test}}$ invariability. The well-trained bundle models predict user-bundle interactions $X$ on $Z_{test}$. The processing of the B-I affiliation is shown in Figure 3. In this way, we can control the degree of B-I affiliaitons variation and simulate the scenario of B-I dynamic variability.
Figure 3. The processing of the B-I affiliation to match the scenario of B-I dynamic variability.
$\bullet$ Hyperparameter Settings We set the hyperparameters of six backbone models from their authors and optimize RDiffBR with the Adam optimizer with a learning rate of 0.001. The time steps $T$ and sampling steps $T^{\prime}$ are set to values ranging from [1,200]. The loss balance factor $\lambda$ is chosen from (0,5), while the residual connection weight $\delta$ is selected from (0,1). The number of MLP layers is selected within 1, 2, 4, 8. The hidden size of the MLP layers are chosen from [8,4096]. Details of other hyperparameters can be found in our released code. All the models are trained using two GTX 3090.
3.2. Overall Performance (RQ1)
Table 2. Performance comparison between backbone models and our RDiffBR on four datasets($\rho$=-4, -2, 0, 2, 4), while $\rho<0$ indicates $Z_{\mathbf{test}}=Z_{\mathbf{train\&val}}\setminus\Delta^{-}$, $\rho>0$ indicates adding set $Z_{\mathbf{test}}=Z_{\mathbf{train\&val}}\cup\Delta^{+}$, and $\rho=0$ indicates $Z_{\mathbf{train\&val}}$ invariability. Improv. indicates the relative improvement of a backbone model+RDiffBR over the itself.

### v3

@@
Abstract
Existing solutions for bundle recommendation (BR) have achieved remarkable effectiveness for predicting the user’s preference for prebuilt bundles. However, bundle-item (B-I) affiliation will vary dynamically in real scenarios. For example, a bundle themed as ‘casual outfit’ may add ‘hat’ or remove ‘watch’ due to factors such as seasonal variations, changes in user preferences or inventory adjustments. Our empirical study demonstrates that the performance of mainstream BR models may fluctuate or decline under item-level variability. This paper makes the first attempt to address the above problem and proposes a novel Residual Diffusion for Bundle Recommendation (RDiffBR) as a model-agnostic generative framework which can assist a BR model in adapting this scenario. During the initial training of the BR model, RDiffBR employs a residual diffusion model to process the item-level bundle embeddings which are generated by the BR model to represent bundle theme via a forward-reverse process. In the inference stage, RDiffBR reverses item-level bundle embeddings obtained by the well-trained bundle model under B-I variability scenarios to generate the effective item-level bundle embeddings. In particular, the residual connection in our residual approximator significantly enhances BR models’ ability to generate high-quality item-level bundle embeddings. Experiments on six BR models and four public datasets from different domains show that RDiffBR improves the performance of Recall and NDCG of backbone BR models by up to 23%, while only increased training time about 4%.
Codehttps://github.com/WUT-IDEA/RDiffBR
Extended versionhttps://arxiv.org/abs/2507.03280
Introduction
@@
Figure 1: An example showing that bundle-item affiliation is changed while bundle theme remains relatively stable.
In practice, platforms can add the ‘hat’ to align with summer promotions or remove the ‘watch’ based on user feedback. An intuitive idea is to regard it as a new bundle. However, such item-level changes may not definitely cause bundle theme shifting although bundle theme is jointly determined by the items within the bundle. As shown in Figure 1, the theme of Bundle 2 and Bundle 3 remain ‘casual outfit’ by item variation. It is advisable to give priority to reusing the identification information of the original bundle because this dynamic variability maintains the theme of bundle.
In recent years, lots of works have focused on bundle recommendation and achieved good results. Including factorization models represented by LIRE (LIRE), EFM (EFM), attention-based models represented by DAM (DAM), Attlist (AttList) and GNN-based models represented by CrossCBR (CrossCBR), BundleGT (BundleGT) and DCBR (DCBR). However, our experiments have found that the performance of existing mainstream bundle recommendation models will fluctuate or even decline under B-I dynamic variability scenarios we have mentioned above. For instance, we found that on Youshu dataset, Recall@20 of BGCN will decrease by up to 19% as items in bundles increases or decreases (see Section 3.2 for details). This is because existing bundle recommendation models lack the ability to effectively model the item-level dynamic variability when generating bundle representations regarding item-level.
The main contributions of this work are summarized as follows. First, we propose a generative residual diffusion framework for bundle recommendation to address item-level dynamic variability issue. Secondly, our residual approximator in diffusion can significantly enhance the item-level bundle embedding generation ability of BR models. Lastly, extensive experiments on four datasets from different domains show that RDiffBR enhances the performance of Recall and NDCG of the BR backbone models by up to 23%, while only increasing the training time by about 4%.
Figure 2: Overview of the proposed RDiffBR framework.
Related Work
Bundle Recommendation. Bundle recommendation aims to recommend a bundle of items related to a specific theme. Early work employed neural network methods for bundle recommendation (EFM; DAM; AttList; LIRE). Recently, graph neural networks have become the mainstream technology for bundle recommendation. We can categorize these graph bundle recommendation models based on these three types of representation learning strategies (Survey_bundle). $\bullet$ Unified learning. The unified representation learning strategy aims to learn representations of users and bundles within a unified user-bundle-item graph (POG; UHBR). $\bullet$ Separate learning. The main idea of this strategy is to first perform representation learning and preference prediction upon two views individually, and then fuse two view-specific predictions to obtain the final prediction (BGCN; IHBR; BundleGT; HBGCN; CMRec). $\bullet$ Cooperative learning. This approach learns user and bundle representations from bundle view and item view and uses techniques such as contrastive learning (MIDGN; CrossCBR; MultiCBR; HIDGN), knowledge distillation (HyperMBR; DGMAE) or diffusion (DCBR) to achieve cross views mutual enhancement. However, most of bundle recommendation lack the ability to adapt to item-level dynamic variability scenario. It is necessary to study our task for real-world application scenarios.
Diffusion Models for Recommendation. The emerging diffusion model in generative models, compared with GANs and VAEs, possess stronger stability and representation. Diffusion model plays three main roles in recommendation systems (survey_diffusion): $\bullet$ Diffusion for data augmentation. This approach generating realistic samples to expand the original training data, such as sequence data (Diff4rec; Plug_diff) and side information (DiffKG; DiffMM). $\bullet$ Diffusion for representation enhancement. Diffusion models converts the input data into embeddings to address some key challenges of recommendation. Lots of studies have been proposed such as denoising the implicit user feedback (DDRM), sequential recommendation (DiffuRec), out-of-distribution issue (CausalDiffRec) and multi-modal learning (MCDRec). $\bullet$ Diffusion as recommendation model. This approach can be classified into three categories: collaborative recommendation (DRM; CFDiff), context-aware recommendation (RecDiff; Diff-POI), and other applications (DCDR). However, a straightforward application of diffusion models on bundle recommendation is insufficient to address the issue of item-level dynamic variability. There are mainly two reasons: the loss targets of diffusion and bundle recommendation are inconsistency and the approximator needs to be designed base on item-level dynamic variability scenario.
Methodology
While it is intuitive to leverage side information—such as item attributes—to enrich the feature space, doing so typically demands non-trivial architectural changes to the original model. In contrast, our RDiffBR operates without any external information and leaves the base architecture untouched. The technical details are presented below.
Problem Definition
$U(|U|=M)$ is denoted as the user set, B(|B|=N)B(|B|=N)as the bundle set, and $I(|I|=O)$ as the item set. The user-bundle interactions, user-item interactions, and bundle-item affiliations are represented as $X_{M\times N}=\{(u,b)\mid u\in U,b\in B\}$, $Y_{M\times O}=\{(u,i)\mid u\in U,i\in I\}$ and $Z_{N\times O}=\{(b,i)\mid b\in B,i\in I\}$ with a binary value at each entry respectively. The value 1 indicates an observed interaction of user-item or user-bundle, or affiliation of bundle-item. The item-level dynamic variability problem can be formulated as follows:
Output: A well-trained bundle recommendation estimates the probability $\hat{y}_{ub}\in(0,1)$ of a user $u$ interacting with unseen bundles $b$ based on $Z^{*}=Z\cup\Delta^{+}\setminus\Delta^{-}$ rather than $Z$. Here, $\Delta^{+}=\{(b,i)\mid(b,i)\notin Z\}$ and $\Delta^{-}=\{(b,i)\mid(b,i)\in Z\}$ represent dynamic variability bundle-item affiliation. And when an item is replaced by others, this can be treated as the joint operation of $\Delta^{+}$ and $\Delta^{-}$.
RDiffBR Framework
Considering the successful experience of diffusion as a generative model and the sparsity of input data, our RDiffBR framework introduces a residual diffusion model in the latent space of interactions. As shown in Figure 2, after BR model utilizes an encoder to get bundle-level user/bundle embeddings $E^{BL}_{u},E^{BL}_{b}$ and item-level embeddings user/bundle embeddings $E^{IL}_{u},E^{IL}_{b}$ from $\{X,Y,Z\}$. Since $E^{IL}_{b}$ represents the bundle theme generated by the BR model, we iteratively process $E^{IL}_{b}$ through a forward-backward process, jointly optimizing the BR model and the residual approximator module that we have designed based on the task. In the inference stage, RDiffBR applies this optimized forward-reverse process to $E^{IL}_{b}$ from dynamic variability B-I affiliations $Z^{*}$ to obtain new $E^{IL}_{b}$, enhancing BR model’s ability for this scenario.
Residual Diffusion
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\boldsymbol{I}).$ (1)
Based on the reparameterization (Diffusion) and the additivity of independent Gaussian noises, we can directly obtain $e_{t}$ from $e_{0}$.
$q(e_{t}|e_{0})=\mathcal{N}(e_{t},\sqrt{\bar{\alpha}_{t}}e_{0},(1-\bar{\alpha_{t}})\boldsymbol{I}),$ (2)
where $\bar{\alpha_{t}}=\prod_{t=1}^{t}\alpha_{t},\alpha_{t}=1-\beta_{t},\beta_{t}\in(0,1)$ controls the Gaussian noise scales added at each step $t$. To regulate the added noises in $e_{1:T}$, we follow  (DRM) using a linear variance noise schedule in the forward process:
$1-\bar{\alpha}_{t}=s\cdot\left[\alpha_{\min}+\frac{t-1}{T-1}(\alpha_{\max}-\alpha_{\min})\right],$ (3)
where T is the total forward step and s $\in$ (0, 1) controls the noise scale. The condition $\alpha_{min}$ << $\alpha_{max}$ $\in$ (0, 1) indicates the upper and lower limits of the added noise.
$p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\Sigma_{\theta}(\hat{e}_{t},t)),$ (4)
where $\hat{e}_{t}$ is denoised embedding in the reverse step t, $\theta$ and are the learnable parameters of residual approximator. This module is executed iteratively in the reverse process until reconstructing final embedding $\hat{e}_{0}$
$q(e_{t-1}|e_{t},e_{0})\propto\mathcal{N}(e_{t-1};\tilde{\mu}(e_{t},e_{0},t),\sigma^{2}(t)\boldsymbol{I}),\mathrm{~where}$ (5)
@@
$\tilde{\mu}(e_{t},e_{0},t)$ and $\sigma^{2}(t)I$ are the mean and covariance of $q(e_{t-1}|e_{t},e_{0})$. Following  (DDRM) we can similarity factorize $p_{\theta}(e_{t-1}|e_{t})$:
@@
$p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\tilde{\beta}_{t}\boldsymbol{I})),\quad\mathrm{~where}$ (6)
@@
$\tilde{e}_{0}=\delta f_{\theta}(\hat{e}_{t},t)+(1-\delta)e_{t^{\prime}},$ (7)
where $\tilde{e}_{0}$ is the predicted value of $e_{0}$ by residual approximator with parameter $\theta$. $\delta$ is residual connection weight. $e_{t^{\prime}}$ is noisy embedding in the forward step $t^{\prime}$. Step information t is encoded by sinusoidal positional encoding (Diffusion), and these two inputs are concated as input.
Loss Function.
@@
$\mathcal{L}_{r}$ $=\mathbb{E}_{q(\mathbf{e}_{t}|\mathbf{e}_{0})}\left[D_{\mathrm{KL}}\left(q\left(\mathbf{e}_{t-1}|\mathbf{e}_{t},\mathbf{e}_{0}\right)\|p_{\theta}\left(\mathbf{e}_{t-1}|\mathbf{e}_{t}\right)\right)\right]$ (8)
@@
Furthermore, by substituting Eq.(5) and Eq.(6) into Eq.(8) and simplifying (DRM) we can drive:
[…]
@@
$\mathcal{L}=\mathcal{L}_{BR\mathrm{~model}}+\lambda\mathcal{L}_{\mathrm{diff}}.$ (10)
@@ 13     Take gradient descent step on $\nabla_{\theta}(\mathcal{L})$ to optimize $\theta$;
Inference.
In the inference stage, we take the item-level bundle embedding $e^{*}_{0}$ in the B-I dynamic variability scenario as input. Subsequently, RDiffBR continuously introduces noise into $e^{*}_{0}$ in the sequence $e^{*}_{0}\to e^{*}_{1}\to...\to e^{*}_{T^{\prime}}$ for $T^{\prime}$ steps in the forward process and then we set $\hat{e}_{{T^{\prime}}}=e^{*}_{T^{\prime}}$ to execute reverse process $\hat{e_{T^{\prime}}}\to\hat{e}_{{T^{\prime}}-1}\to...\to\hat{e}_{0}$ to generate new item-level bundle embedding $\hat{e}_{0}$. The reverse ignores the variance and uses $\hat{e}_{t-1}=\mu_{\theta}(\hat{e}_{t},e^{*}_{0},t)$ for directly inference following (MultiVAE). We replace the original $e^{*}_{0}$ with $\hat{e}_{0}$, and then conduct bundle recommendation following the original way of the bundle model. Finally the ranking of bundles is enhanced. The inference procedure of residual approximator is illustrated in Algorithm 2. Specifically, Line 1 represent the forward process, Lines 2-5 represent the reverse process using a residual approximator to obtain ideal item-level bundle embedding.
@@
Experiments
Our experiments aim to answer the following questions:
$\bullet$ RQ1: How does RDiffBR perform with backbone bundle models under various degrees of B-I variability?
$\bullet$ RQ2: What is the contribution of residual connection in RDiffBR to its overall performance?
@@
Experimental Settings
$\bullet$ Datasets. Four publicly bundle datasets from different domains are summarized in Table 1. Youshu (DAM) for book list recommendation; NetEase (EFM) for music playlist recommendation; iFashion (POG) for fashion outfit recommendation; In addition, we introduce another meal recommendation dataset $\mathrm{MealRec+}$ (mealrec+), where the meal consisted of individual course items is treated as bundle.
[…]
# U 8,039 18,528 53,897 1,575
# I 32,770 123,628 42,563 7,280
# B 4,771 22,864 27,694 3,817
# U-I 138,515 1,128,065 2,290,645 151,148
# U-B 51,377 302,303 1,679,708 46,767
# B-I 176,667 1,778,838 106,916 11,451
# Avg. I/B 37.03 77.8 3.86 3
Dens.B-I 0.11% 0.06% 0.01% 0.77%
Table 1: Statistics of datasets.